I'm stuck on the below:
t = u + uqp ;solve for u
I divide both sides by qp but cant figure out what to do about the u + u...
Please assist

Answer by Alan3354(69443)

t = u + uqp ;solve for u
------
t = u*(1 + qp)
u = t/(1 + qp)